OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6.

OpenStudy (♪chibiterasu):

Options?

OpenStudy (anonymous):

y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Ibbutibbu.

OpenStudy (ibbutibbu.):

yes i can help you let me se...

OpenStudy (ibbutibbu.):

A is the answer

OpenStudy (ibbutibbu.):

do you need an explanation